Parth (parthkohli):

\[\large \rm \lim_{x \to 0}{\sin(3x) \over x}\]

Parth (parthkohli):

I have a feeling that we have to use\[\rm \large\lim_{x \to 0}{\sin(x) \over x} = 1\]

OpenStudy (anonymous):

Squeeze theorem?

OpenStudy (raden):

just multiply by 3/3

Parth (parthkohli):

How about multiplying 3 to both numerator and denominator?\[\large \rm\lim_{x \to 0}{\sin(3x) \over 3x}\times 3\]Oh

Parth (parthkohli):

Hehe, thanks!

OpenStudy (raden):

welcome :)

OpenStudy (anonymous):

Haha, nice hack. ;)